OpenStudy (ttop0816):

Please help! Evaluate tan(-x/6). -squareroot3/3 -squareroot 3 -1

OpenStudy (agent0smith):

Er, are you sure that's written correctly?

OpenStudy (agent0smith):

Guessing it's this. pi/6 not x/6.\[\Large \tan \left( - \frac{ \pi }{ 6 }\right)\]

OpenStudy (ttop0816):

@agent0smith yes! im sorry! my mistake

OpenStudy (kc_kennylau):



OpenStudy (kc_kennylau):

pi/6 is just 60deg

OpenStudy (agent0smith):

@ttop0816 you should have notes or a book which gives the value of tan(pi/6). Then from that you decide the value of tan(-pi/6), since tangent is negative in the fourth quadrant.

OpenStudy (ttop0816):

then would the answer be the second one?

OpenStudy (agent0smith):

What second one?

OpenStudy (ttop0816):

-squareroot 3...?

OpenStudy (agent0smith):

Well what is the value of tan(pi/6)? You shouldn't have to guess on that one, use your book or notes and there should be a table.

OpenStudy (ttop0816):

oh its -square root 3/3 then!

OpenStudy (agent0smith):

Hopefully you've seen this before